22west:

Add the two expressions. 3z−4 and 2z + 5

22west:

@dude

dude:

We have to combine like terms 3z-4+2z+5

22west:

3z and 2z 4 and 5

dude:



22west:

ok first of all the circles and lines look like sperm cells

dude:

Lmao thanks

22west:

just sayin

22west:

pun intended

dude:

Lmao, back to the question, combine the terms When adding, remember to take the sign of the bigger number

22west:

umm

dude:

-4+5 3z+2z

22west:

is that the answer

dude:

No, I am showing you what to add

22west:

oh ok

22west:

-4 + 5z?

dude:

What is -4+5?

dude:

Or 5-4

22west:

1

dude:

Yes, so it is 1+5z

22west:

Add the two expressions. m + 6 and 4m + 2

dude:

add 4m and m add 6 and 2

22west:

4m + m 8

dude:

m is just 1m So 4m +1m

22west:

oh

22west:

5m

dude:

Yes so it is 5m+8

22west:

Add the two expressions. −2.4n−3 and −7.8n+2

dude:

-2.4n+-7.8n= -3+2=

22west:

-3+2=-1 -2.4+(-7.8)

22west:

=10.2

dude:

Keep the sign!

22west:

-

dude:

-10.2n-1

22west:

that is the answer?

dude:

Yes

22west:

Subtract m + 8 ​​from​ 5m + 11.

dude:

\((5m+11)-(m+8)\) Distribute the -1 into the second set of parenthesis

22west:

(5m+11)-(1m+8)

22west:

thats the answer?

dude:

No, you have to distribute -1 into (m+8)